OpenStudy (anonymous):

What is the y-value for f(x) = cos(x) when x = -90°?

OpenStudy (anonymous):

@demonchild99 @william.1233 @YGHOOTIE @dejmccray @haydenhd @Vendela

OpenStudy (anonymous):

51

OpenStudy (anonymous):

Put \(x=-90\)

OpenStudy (anonymous):

\[f(x) = \cos(-90^{\circ})\]

OpenStudy (anonymous):

And remember, cos(x) is an even function that means: \[\cos(-x) = \cos(x)\] It means negative angle just becomes positive..

OpenStudy (anonymous):

so the answere will be? :D

OpenStudy (anonymous):

\[f(x) = \cos(-90^{\circ}) \\ f(x) = \cos(90^{\circ})\] Find out what is cos(90), and f(x) is nothing by \(y-value\)..

OpenStudy (anonymous):

*but

OpenStudy (anonymous):

but? @eta?

OpenStudy (anonymous):

ok never mind i see it now lol thank u so very much

OpenStudy (anonymous):

but you only want answer, you do not want to learn the things.. :P

OpenStudy (anonymous):

I wrote "nothing by y-value, there is no by, by should be but there..

OpenStudy (anonymous):

never said that wat u told me actually help me figure out my last 5 questions i have

OpenStudy (anonymous):

-.- not trying to be mean but please dont under dtimate people @eta

OpenStudy (anonymous):

Hey, do I still need to mention that I was just kidding? I told my expression :P with that statement..

OpenStudy (anonymous):

Okay, allow me to ask for sorry, I am sorry if you felt any bad from my side.. :(